A metal particles travels down ward through a fluid that extend from plate A and plate B under the influence of magnetic field. If the particles is released from test at mid point. The distance= 100m and acceleration 45m/s. Determine the velocity when it reaches plate B.



1
Expert's answer
2021-03-12T07:15:20-0500

Acceleration(a) = 45 m/s2

Distance(s) = 100 m

Initial velocity(u) = 0 m/s

From Newton's laws of motion,

"v^2-u^2=2as\\\Rightarrow v=\sqrt{2as}\\\Rightarrow v=\sqrt{2\times45\times100}\\\Rightarrow v=94.86\space m/s"